# Feature Re-Embedding: Towards Foundation Model-Level Performance in Computational Pathology

### 27 Feb 2024 v1

@@ 3.2 Re-embedded Regional Transformer
where is the mapping function of the RTransformer here, is the processed input instance features, and is the new embedding of the -th instance. is the dimensionality of the embedding. The RTransformer can be flexibly plugged into the MIL framework as a re-embedding module after feature input and before instance aggregation to reduce bias caused by the shift between offline feature learning and downstream tasks. The whole re-embedding process in the RTransformer can be formulated as,
@@
| ResNet-50 ImageNet pretrained | AB-MIL [14] | 90.060.72 | 94.540.30 | 87.830.83 | 85.742.77 | 91.402.41 | 71.314.23 |
| CLAM [22] | 90.140.85 | 94.700.76 | 88.100.63 | 85.172.70 | 91.671.78 | 70.684.15 | |
| DSMIL [18] | 90.171.02 | 94.570.40 | 87.651.18 | 87.202.69 | 91.581.33 | 73.263.98 | |
| TransMIL [26] | 89.222.32 | 93.512.13 | 85.104.33 | 85.153.13 | 91.381.93 | 70.903.81 | |
| DTFD-MIL [45] | 90.220.36 | 95.150.14 | 87.620.59 | 88.332.05 | 91.551.01 | 73.502.88 | |
| IBMIL [20] | 91.230.41 | 94.801.03 | 88.800.89 | 88.253.25 | 91.352.09 | 74.494.34 | |
| MHIM-MIL [30] | 91.810.82 | 96.140.52 | 89.940.70 | 83.903.38 | 92.481.38 | 70.003.76 | |
| RT-MIL | 92.400.31 | 97.320.29 | 90.630.45 | 88.330.67 | 93.171.45 | 75.021.58 | |
| PLIP WSI pretrained | AB-MIL [14] | 94.660.42 | 97.300.31 | 93.290.54 | 86.022.88 | 91.801.81 | 71.313.79 |
| CLAM [22] | 93.730.54 | 97.170.50 | 91.600.60 | 86.701.35 | 92.162.02 | 72.612.71 | |
| DSMIL [18] | 94.400.85 | 97.060.56 | 92.781.15 | 87.252.70 | 91.801.67 | 72.732.49 | |
| TransMIL [26] | 94.400.43 | 97.880.21 | 92.810.43 | 87.094.02 | 90.852.76 | 72.756.39 | |
| DTFD-MIL [45] | 94.570.31 | 97.290.23 | 93.120.40 | 85.443.27 | 91.782.44 | 70.625.04 | |
| IBMIL [20] | 93.900.66 | 97.040.18 | 92.440.64 | 86.500.44 | 91.581.95 | 72.330.66 | |
| MHIM-MIL [30] | 95.320.31 | 97.790.15 | 94.130.42 | 87.661.24 | 92.702.05 | 74.031.59 | |
| RT-MIL | 95.490.00 | 98.050.29 | 94.290.04 | 87.852.36 | 93.711.11 | 75.003.40 | |
@@ 4.3.1 Re-Embedding Matters in Pathology
Different Re-embedding Discussion: The bottom part of Table 4 presents the performance of AB-MIL under different feature settings. We employ four methods, including TranMIL, N-MSA, the local version of N-MSA, and our proposed RT to re-embed the features. Notably, we are the first to apply the existing Transformer for instance feature re-embedding, which was not mentioned at all in the original literature. From observations, all four re-embedded AB-MILs perform better than the original one. The performance improvement by TransMIL, N-MSA, and RT are 1.28%, 1.68%, and 2.80%, respectively, on C16, while these numbers on LUAD are 4.46%, 4.14%, and 8.41%, respectively. This phenomenon validates the importance of re-embedding in MIL-based computational pathology. Among the four employed re-embedding approaches, RT boosts the AB-MIL with the most considerable improvements while incurring the lowest computational cost. Specifically, RT achieves more performance gains (+0.85% on C16 and +2.63% on LUAD) while only requiring a 1/5 inference time compared with the second-best approaches.
@@ 4.3.4 Impact of Cross-Region MSA
![[Uncaptioned image]](extracted/5433878/imgs/cw_msa.png)
@@ 5 Conclusion
In this work, we have demonstrated the importance of instance feature re-embedding for computational pathology algorithms based on MIL, alleviating the issue of the under-learning of instance features in the conventional MIL paradigm. We have also shown that Transformer-based re-embedding modules can consistently boost the performance of various MIL methods regardless of their architectures. However, the main result of this paper is the introduction of the Re-embedded Regional Transformer and two novel components: CR-MSA and EPEG. We have evidence of the importance of the local Transformer in the age of the foundation model and its versatility as a re-embedding module. In the future, we plan to explore how feature re-embedding can facilitate advanced tasks such as localization and segmentation in computational pathology.
@@ D.4 More Parameters are not Always Better
Generally, a Transformer is a multi-layer structure that stacks several blocks with the same structure [21, 33, 17, 9]. However, due to the task specificity, RT only contains a few blocks. Here, we systematically investigate the impact of different numbers of layers and different blocks on the performance and computational cost of RT. First, Figure 9 shows two different blocks: (a) is used by default in RT; (b) introduces a feed-forward network (FFN), which plays an indispensable role in Transformer for NLP or natural image computer vision tasks. From Table 8, we can find that FFN introduces a large number of parameters and computation, but the more expensive computation cost does not bring performance improvement. Similarly, increasing the number of layers comes with nearly twice the computational cost, but the performance drops significantly. Our experiments demonstrate that lightweight Transformer shows superiority in performance and efficiency for computational pathology.
| Model | C16 | NSCLC | LUAD | Para. |
| [w/o FFN] | 97.32 | 96.40 | 67.19 | 2.70M |
| 96.23 | 95.58 | 64.89 | 6.90M | |
| 96.44 | 96.31 | 65.15 | 3.75M | |
| 96.57 | 95.70 | 63.07 | 10.05M |

### 07 Apr 2024 v2

@@ 3.2 Re-embedded Regional Transformer
where is the mapping function of the RTransformer here, is the processed input instance features, and is the new embedding of the -th instance. is the dimension of the embedding. The RTransformer can be flexibly plugged into the MIL framework as a re-embedding module after feature input and before instance aggregation to reduce bias caused by the shift between offline feature learning and downstream tasks. The whole re-embedding process in the RTransformer can be formulated as,
@@
| ResNet-50 ImageNet pretrained | AB-MIL [14] | 90.060.72 | 94.540.30 | 87.830.83 | 86.414.92 | 91.102.52 | 81.644.71 |
| CLAM [22] | 90.140.85 | 94.700.76 | 88.100.63 | 85.172.70 | 91.671.78 | 80.373.04 | |
| DSMIL [18] | 90.171.02 | 94.570.40 | 87.651.18 | 87.202.69 | 91.581.33 | 82.412.92 | |
| TransMIL [26] | 89.222.32 | 93.512.13 | 85.104.33 | 84.682.67 | 90.801.91 | 79.862.63 | |
| DTFD-MIL [45] | 90.220.36 | 95.150.14 | 87.620.59 | 85.921.76 | 91.431.64 | 81.092.05 | |
| IBMIL [20] | 91.230.41 | 94.801.03 | 88.800.89 | 84.193.40 | 91.012.32 | 79.453.42 | |
| MHIM-MIL [30] | 91.810.82 | 96.140.52 | 89.940.70 | 86.735.59 | 92.361.58 | 82.435.47 | |
| RT-MIL | 92.400.31 | 97.320.29 | 90.630.45 | 88.330.67 | 93.171.45 | 83.700.95 | |
| PLIP WSI pretrained | AB-MIL [14] | 94.660.42 | 97.300.31 | 93.290.54 | 85.452.32 | 91.732.26 | 80.602.66 |
| CLAM [22] | 93.730.54 | 97.170.50 | 91.600.60 | 86.701.35 | 92.162.02 | 81.911.78 | |
| DSMIL [18] | 94.400.85 | 97.060.56 | 92.781.15 | 87.252.70 | 91.801.67 | 82.182.28 | |
| TransMIL [26] | 94.400.43 | 97.880.21 | 92.810.43 | 85.833.44 | 92.172.20 | 81.123.25 | |
| DTFD-MIL [45] | 94.570.31 | 97.290.23 | 93.120.40 | 86.422.67 | 92.162.42 | 81.772.73 | |
| IBMIL [20] | 93.900.66 | 97.040.18 | 92.440.64 | 87.571.48 | 91.711.74 | 82.782.02 | |
| MHIM-MIL [30] | 95.320.31 | 97.790.15 | 94.130.42 | 87.072.20 | 93.172.00 | 82.482.50 | |
| RT-MIL | 95.490.00 | 98.050.29 | 94.290.04 | 88.823.22 | 93.801.24 | 84.553.55 | |
@@ 4.3.1 Re-Embedding Matters in Pathology
Different Re-embedding Discussion: The bottom part of Table 4 presents the performance of AB-MIL under different feature settings. We employ four methods, including TranMIL, N-MSA, the local version of N-MSA, and our proposed RT to re-embed the features. From observations, all four re-embedded AB-MILs perform better than the original one. The performance improvement by TransMIL, N-MSA, and RT are 1.28%, 1.68%, and 2.80%, respectively, on C16, while these numbers on LUAD are 4.46%, 4.14%, and 8.41%, respectively. This phenomenon validates the importance of re-embedding in MIL-based computational pathology. Among the four employed re-embedding approaches, RT boosts the AB-MIL with the most considerable improvements while incurring the lowest computational cost. Specifically, RT achieves more performance gains (+0.85% on C16 and +2.63% on LUAD) while only requiring a 1/5 inference time compared with the second-best approaches.
@@ 4.3.4 Impact of Cross-Region MSA
![[Uncaptioned image]](extracted/5521055/imgs/cw_msa.png)
@@ 5 Conclusion
In this work, we have demonstrated the importance of instance feature re-embedding for computational pathology algorithms based on MIL, alleviating the issue of the under-learning of instance features in the conventional MIL paradigm. We have also shown that Transformer-based re-embedding modules can consistently boost the performance of various MIL methods regardless of their architectures. However, the main result of this paper is the introduction of the Re-embedded Regional Transformer and two novel components: CR-MSA and EPEG. We have evidence of the importance of the local Transformer in the age of the foundation model and its versatility as a re-embedding module.
@@ D.3 Region Partition Strategy
| PretrainingData | Para. | TrainTime | Memo. | FPS | C16 | NSCLC | LUAD | |
| ABMIL | ImageNet-1K | 0.65M | 3.1s | 2.3G | 1250 | 94.54 | 95.28 | 58.78 |
| ABMIL+PLIP [11] | OpenPath-200K | 0.65M | 1.7s | 2.2G | 2273 | 97.30 | 95.68 | 62.09 |
| DTFD [41] | ImageNet-1K | 0.79M | 5.1s | 2.1G | 325 | 95.15 | 95.55 | 59.48 |
| TransMIL [22] | ImageNet-1K | 2.67M | 13.2s | 10.6G | 76 | 93.51 | 94.97 | 64.11 |
| Re-embedding | ||||||||
| ABMIL+N-MSA [36] | ImageNet-1K | 1.64M | 7.7s | 7.2G | 158 | 96.20 | 95.51 | 63.99 |
| RT-MIL(w/o CR-MSA) | ImageNet-1K | 1.64M | 6.1s | 10.0G | 272 | 96.89 | 96.24 | 63.03 |
| RT-MIL | ImageNet-1K | 2.70M | 6.5s | 10.1G | 236 | 97.32 | 96.40 | 67.19 |
| RT-MIL [w/ FFN] | ImageNet-1K | 6.90M | 9.9s | 12.0G | 114 | 96.23 | 95.58 | 64.89 |
| RT-MIL [w/ FFN] | ImageNet-1K | 10.05M | 14.4s | 16.3G | 71 | 96.57 | 95.70 | 63.07 |
@@ D.4 More Parameters are not Always Better
Generally, a Transformer is a multi-layer structure that stacks several blocks with the same structure [21, 33, 17, 9]. However, due to the task specificity, RT only contains a few blocks. Here, we systematically investigate the impact of different numbers of layers and different blocks on the performance and computational cost of RT. First, Figure 9 shows two different blocks: (a) is used by default in RT; (b) introduces a feed-forward network (FFN), which plays an indispensable role in Transformer for NLP or natural image computer vision tasks. From Table 8, we can find that FFN introduces a large number of parameters and computation, but the more expensive computation cost does not bring performance improvement.
Moreover, we can summarize that: 1) Transformer-based methods, represented by TransMIL, bring better long-sequence modeling ability, but also introduce several times more parameters. RT-MIL, as one of them, has equal or less parameter size compared to TransMIL and better performance. 2) As Transformer-based methods, Re-embedding paradigm and RT have Higher parameter efficiency. “+N-MSA” has the same model structure as TransMIL (N-MSA2), but thanks to the Re-embedding, it can achieve higher performance with lower cost. Moreover, RT (w/o CR-MSA) leverages a more excellent design to further improve performance and training time. This indicates that RT-MIL has a good parameter compression space, and can achieve significant improvement with 2 parameters compared to DTFD, reaching the level of the foundation model. 3) In computational pathology, limited by the number of slides, the models face the problem of over-fitting[45], and higher parameter size does not imply better performance. Not only does TransMIL perform poorly on some tasks, we also add FFN and increase the number of layers on the RT-MIL to increase the parameters of RT-MIL (+7.35M), but the performance drops significantly (-4.12% on LUAD).